# Partial Rewriting and Value Interpretation of Logically Constrained Terms

### v1

@@ 2 Preliminaries
The constraint rewrite rules in Example˜1.1 in our formulation are presented as $\Pi\{\hskip 0.85358ptx\hskip 0.85358pt\}.\ \mathsf{sum}(x)\rightarrow 0\penalty 10000\ [\hskip 0.85358pt0\geqslant x\hskip 0.85358pt]$ and $\Pi\{\hskip 0.85358ptx\hskip 0.85358pt\}.\ \mathsf{sum}(x)\rightarrow x+\mathsf{sum}(x+-1)\penalty 10000\ [\hskip 0.85358ptx>0\hskip 0.85358pt]$, both of which are left-linear. Let $s$ be a constrained term $\Pi\{\hskip 0.85358ptx\hskip 0.85358pt\}.\ \mathsf{sum}(x)\penalty 10000\ [\hskip 0.85358ptx>2\hskip 0.85358pt]$. To perform a rewrite step from $s$, we take a renamed variant of the rule such as $\rho:\Pi\{\hskip 0.85358ptx^{\prime}\hskip 0.85358pt\}.\ \mathsf{sum}(x^{\prime})\rightarrow x^{\prime}+\mathsf{sum}(x^{\prime}+-1)\penalty 10000\ [\hskip 0.85358ptx^{\prime}>0\hskip 0.85358pt]$. Then $s$ has a $\rho$-redex at position $\epsilon$ using the substitution $\gamma=\{x^{\prime}\mapsto x\}$, as $\vDash_{\mathcal{M}}x>2\Rightarrow x>0$. From this we obtain a most general rewrite step $s\to^{\epsilon}_{\rho,\gamma}\Pi\{x\}.\ x+\mathsf{sum}(x+-1)\penalty 10000\ [\hskip 0.85358ptx>2\land x>0\hskip 0.85358pt]$. In this case, the additional constraint $x>0$ is auxiliary, but for example, for the rewrite rule $\rho^{\prime}:\Pi\{\hskip 0.85358ptx^{\prime},y^{\prime}\hskip 0.85358pt\}.\ \mathsf{f}(x^{\prime})\rightarrow x^{\prime}+y^{\prime}\penalty 10000\ [\hskip 0.85358pty^{\prime}\geqslant x^{\prime}\hskip 0.85358pt]$, this addition is necessary, e.g. $\Pi\{x\}.\ \mathsf{f}(x)\penalty 10000\ [\hskip 0.85358ptx>2\hskip 0.85358pt]\to_{\rho^{\prime}}\Pi\{y^{\prime}\}.\ x+y^{\prime}\penalty 10000\ [\hskip 0.85358ptx>2\land y^{\prime}>x\hskip 0.85358pt]$. Note that the extra variable $y^{\prime}$ in the rule remains not instantiated as the domain of the substitution equals to $\mathcal{V}\mathsf{ar}(\ell)$.
@@ 3 Partial Reduction of Constrained Terms
Let $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ be a satisfiable existentially constrained term. Suppose that $\rho\colon\Pi Z.\ \ell\rightarrow r\penalty 10000\ [\hskip 0.85358pt\pi\hskip 0.85358pt]$ is a left-linear constrained rule satisfying $\mathcal{V}\mathsf{ar}(\rho)\cap\mathcal{V}\mathsf{ar}(s,\varphi)=\varnothing$. We say $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ has a partial $\rho$-redex at position $p\in\mathcal{P}\mathsf{os}(s)$ using substitution $\gamma$ if the following is satisfied: 1. $\mathcal{D}\mathsf{om}(\gamma)=\mathcal{V}\mathsf{ar}(\ell)$, 2. $s|_{p}=\ell\gamma$, 3. $\gamma(x)\in\mathcal{V}\mathsf{al}\cup X$ for all $x\in\mathcal{V}\mathsf{ar}(\ell)\cap Z$, and 4. $(\exists\vec{x}.\ \varphi)\land(\exists\vec{z}.\ \pi\gamma)$ is satisfiable in $\mathcal{M}$ where $\{\hskip 0.85358pt\vec{z}\hskip 0.85358pt\}=\mathcal{V}\mathsf{ar}(\pi)\setminus\mathcal{V}\mathsf{ar}(\ell)$.
@@
We now compare partial rewrite steps with the most general rewrite steps of existentially constrained terms. First, we show that every most general rewrite step is also a partial rewrite step.
@@
Proof 3.7.
As we demonstrated in Example˜3.5, partial rewrite steps may not be most general rewrite steps. Thus, the converse does not hold in general. However, we present a characterization of the converse.
@@
Proof 3.9.
By $\{\hskip 0.85358pt\vec{z}\hskip 0.85358pt\}=\mathcal{V}\mathsf{ar}(\pi)\setminus\mathcal{V}\mathsf{ar}(\ell)$ and condition (1) we have $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{z}.\ \pi\gamma)\subseteq\{\gamma(x)\mid x\in\mathcal{V}\mathsf{ar}(\ell)\cap\mathcal{V}\mathsf{ar}(\pi)\}$. Then, since $\mathcal{V}\mathsf{ar}(\pi)\subseteq Z$ by the definition of rewrite rule, it follows that (4.5) $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{z}.\ \pi\gamma)\subseteq X$ by condition (3).
The following is implied by LABEL:lem:from_weak_rewrite_step_to_mostgeneral_step.

### v2

@@ 2 Preliminaries
For the LCTRS in Example˜1.1 we have $\mathcal{S}_{\mathsf{th}}=\{\hskip 0.85358pt\mathsf{Bool},\mathsf{Int}\hskip 0.85358pt\}$, $\mathcal{S}_{\mathsf{te}}=\varnothing$, $\mathcal{F}_{\mathsf{th}}=\mathbb{Z}\cup\mathbb{B}\cup\{\hskip 0.85358pt+:\mathsf{Int}\times\mathsf{Int}\to\mathsf{Int},{\geqslant}:\mathsf{Int}\times\mathsf{Int}\to\mathsf{Bool},{<}:\mathsf{Int}\times\mathsf{Int}\to\mathsf{Bool},\ldots\hskip 0.85358pt\}$, and $\mathcal{F}_{\mathsf{te}}=\{\hskip 0.85358pt\mathsf{sum}:\mathsf{Int}\times\mathsf{Int}\to\mathsf{Int}\hskip 0.85358pt\}$, where $\mathcal{V}\mathsf{al}=\mathbb{Z}\cup\mathbb{B}$. We have $\mathcal{I}(\mathsf{Bool})=\mathbb{B}$, $\mathcal{I}(\mathsf{Int})=\mathbb{Z}$, $\mathcal{J}(+):\mathbb{Z}\times\mathbb{Z}\to\mathbb{Z}$ is the addition function on $\mathbb{Z}$, and so on. If we take a valuation $\rho$ such that $\rho(x)=\rho(y)=1$, then we have ${x+y}]\!]_{\mathcal{M},\rho}=2$ and ${x+y\geqslant 2}]\!]_{\mathcal{M},\rho}=\mathsf{true}$. Thus, $\vDash_{\mathcal{M},\rho}x+y\geqslant 2$, as $x+y\geqslant 2$ is a constraint. The constraint $x+y\geqslant 2$ is satisfiable but not valid. For example, $x=_{\mathbb{Z}}x$, $(x\geqslant 0)\lor(x<0)$ and $(0<x)\Rightarrow(x\geqslant 0)$ are valid constraints, hence we obtain $\vDash_{\mathcal{M}}(x\geqslant 0)\lor(x<0)$. Note that values are terms, therefore $\sigma=\{\hskip 0.85358ptx\mapsto 1,y\mapsto 1,z\mapsto\mathsf{sum}(w)\hskip 0.85358pt\}$ is a substitution. We have $(x+z)\sigma=1+\mathsf{sum}(w)$. As $\mathcal{V}\mathcal{D}\mathsf{om}(\sigma)=\{\hskip 0.85358ptx,y\hskip 0.85358pt\}$, the substitution $\sigma$ is $\{\hskip 0.85358ptx,y\hskip 0.85358pt\}$-valued which implies $\sigma\vDash_{\mathcal{M}}x+y\geqslant 2$.
An existential constraint is a pair $\langle\vec{x},\varphi\rangle$ of a sequence of variables $\vec{x}$ and a constraint $\varphi$, written as $\exists\vec{x}.\ \varphi$, such that $\{\hskip 0.85358pt\vec{x}\hskip 0.85358pt\}\subseteq\mathcal{V}\mathsf{ar}(\varphi)$. We define the sets of free and bound variables of $\exists\vec{x}.\ \varphi$ as follows: $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{x}.\ \varphi)=\mathcal{V}\mathsf{ar}(\varphi)\setminus\{\hskip 0.85358pt\vec{x}\hskip 0.85358pt\}$ and $\mathcal{B}\mathcal{V}\mathsf{ar}(\exists\vec{x}.\ \varphi)=\{\hskip 0.85358pt\vec{x}\hskip 0.85358pt\}$. We define the interpretation of existential constraints as follows: We have $\vDash_{\mathcal{M},\rho}\exists\vec{x}.\ \varphi$ if there exists a valuation $\rho^{\prime}=\rho[\vec{x}\mapsto\vec{w}]$, i.e. $\rho^{\prime}(x)=\rho(x)$ for any $x\in\mathcal{V}\setminus\{\hskip 0.85358pt\vec{x}\hskip 0.85358pt\}$ and $\rho^{\prime}(\vec{x})=\vec{w}$, such that $\mathcal{M}\vDash_{\mathcal{M},\rho^{\prime}}\varphi$. An existentially constrained term is a triple $\langle X,s,\exists\vec{x}.\ \varphi\rangle$, written as $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$, of a set $X$ of variables, a term $s$, and an existentially constraint $\exists\vec{x}.\ \varphi$, such that $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{x}.\ \varphi)\subseteq X\subseteq\mathcal{V}\mathsf{ar}(s)$, and $\mathcal{B}\mathcal{V}\mathsf{ar}(\exists\vec{x}.\ \varphi)\cap\mathcal{V}\mathsf{ar}(s)=\varnothing$. Variables in $X$ are called logical variables (of $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$). An existentially constrained term $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ is said to be satisfiable if $\exists\vec{x}.\ \varphi$ is satisfiable.
Example 2.2.
Let us continue with the LCTRS and the sorted signature in Examples˜1.1 and 2.1. Then $\exists x.\ x+y\geqslant 2$ is an example of an existential constraint for which we have $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists x.\ x+y\geqslant 2)=\{\hskip 0.85358pty\hskip 0.85358pt\}$ and $\mathcal{B}\mathcal{V}\mathsf{ar}(\exists x.\ x+y\geqslant 2)=\{\hskip 0.85358ptx\hskip 0.85358pt\}$. Also $\exists x,y.\ x+y\geqslant 2$ and $x+y\geqslant 2$ are existential constraints. Note that for the latter we drop the sequence of variables as it is empty. On the other hand, we exclude $\exists z.\ x+y\geqslant 2$ from the set of existential constraints, as $z\notin\mathcal{V}\mathsf{ar}({x+y\geqslant 2})$. Consider $\rho$ from the example above that yields $\vDash_{\mathcal{M},\rho}\exists w.\ x+y\geqslant w$. Existential constraints can be used to construct the following two examples of existentially constrained terms $\Pi\{\hskip 0.85358ptx,y\hskip 0.85358pt\}.\ \mathsf{sum}(x,y)\penalty 10000\ [\hskip 0.85358pt\exists w.\ x+y\geqslant w\hskip 0.85358pt]$ and $\Pi\{\hskip 0.85358ptx,y\hskip 0.85358pt\}.\ \mathsf{sum}(x,y)\penalty 10000\ [\hskip 0.85358ptx\geqslant y\hskip 0.85358pt]$. The $\Pi$-notation is designed to allow value-only instantiations for variables that do not appear in the constraint part. For example, the constrained term $\Pi\{\hskip 0.85358ptx\hskip 0.85358pt\}.\ \mathsf{sum}(x,y)\penalty 10000\ [\hskip 0.85358pt\mathsf{true}\hskip 0.85358pt]$ allows the instantiations $n\in\mathbb{Z}$ and $t\in\mathcal{T}(\mathcal{F},\mathcal{V})$ in the term $\mathsf{sum}(n,t)$.
An existentially constrained term $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ is said to be subsumed by an existentially constrained term $\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$, denoted by $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\mathrel{\begin{subarray}{c}\textstyle\subset\\[-0.86108pt] \textstyle\sim\end{subarray}}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$, if for all $X$-valued substitutions $\sigma$ with $\sigma\vDash_{\mathcal{M}}\exists\vec{x}.\ \varphi$ there exists a $Y$-valued substitution $\gamma$ with $\gamma\vDash_{\mathcal{M}}\exists\vec{y}.\ \psi$ such that $s\sigma=t\gamma$. Two existentially constrained terms $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ and $\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$ are said to be equivalent, denoted by $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\sim\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$ if $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\mathrel{\begin{subarray}{c}\textstyle\subset\\[-0.86108pt] \textstyle\sim\end{subarray}}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$ and $\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]\mathrel{\begin{subarray}{c}\textstyle\subset\\[-0.86108pt] \textstyle\sim\end{subarray}}\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$.
Example 2.3.
Let us continue with Example˜2.2. We obtain that $\Pi\{\hskip 0.85358ptx,y\hskip 0.85358pt\}.\ \mathsf{sum}(x,y)\penalty 10000\ [\hskip 0.85358ptx\geqslant y\hskip 0.85358pt]\mathrel{\begin{subarray}{c}\textstyle\subset\\[-0.86108pt] \textstyle\sim\end{subarray}}\Pi\{\hskip 0.85358ptx^{\prime},y^{\prime}\hskip 0.85358pt\}.\ \mathsf{sum}(x^{\prime},y^{\prime})\penalty 10000\ [\hskip 0.85358pt\exists w.\ w\geqslant-1\land x^{\prime}=y^{\prime}+w\hskip 0.85358pt]$ and $\Pi\{\hskip 0.85358ptx,y\hskip 0.85358pt\}.\ \mathsf{sum}(x,y)\penalty 10000\ [\hskip 0.85358pt\mathsf{true}\hskip 0.85358pt]\mathrel{\begin{subarray}{c}\textstyle\subset\\[-0.86108pt] \textstyle\sim\end{subarray}}\Pi\{\hskip 0.85358ptx^{\prime}\hskip 0.85358pt\}.\ \mathsf{sum}(x^{\prime},y^{\prime})\penalty 10000\ [\hskip 0.85358pt\mathsf{true}\hskip 0.85358pt]$, but both do not satisfy equivalence. We have, for example, $\Pi\{\hskip 0.85358ptx,y\hskip 0.85358pt\}.\ \mathsf{sum}(x,y)\penalty 10000\ [\hskip 0.85358ptx=1\land y>x\hskip 0.85358pt]\sim\Pi\{\hskip 0.85358pty^{\prime}\hskip 0.85358pt\}.\ \mathsf{sum}(1,y^{\prime})\penalty 10000\ [\hskip 0.85358pty^{\prime}\geqslant 2\hskip 0.85358pt]$. For various characterizations of equivalence, see [17].
A constrained rewrite rule is a quadruple consisting of a set $Z$ of theory variables, terms $\ell,r$ of a same sort and a constraint $\pi$ such that $\mathcal{V}\mathsf{ar}(\pi)\cup(\mathcal{V}\mathsf{ar}(r)\setminus\mathcal{V}\mathsf{ar}(\ell))\subseteq Z$, which is written as $\Pi Z.\ \ell\rightarrow r\penalty 10000\ [\hskip 0.85358pt\pi\hskip 0.85358pt]$. A constrained rule $\Pi Z.\ \ell\rightarrow r\penalty 10000\ [\hskip 0.85358pt\pi\hskip 0.85358pt]$ is said to be left-linear (left-value-free) if $\ell$ is linear (value-free). The original definition of constrained rewriting in [8, 4] is defined modulo equivalence $\sim$ of constrained terms, which makes analyses more complicated and its implementation mostly infeasible. However, most general constrained rewriting on existential constrained terms [18] has several advantages compared to the original definition of constrained rewriting, even though it is presently formulated only for left-linear constrained rewrite rules. Since this paper follows the line of our previous work [18], we again restrict our attention to left-linear constrained rewrite rules.
Let $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ be a satisfiable existentially constrained term. Suppose that $\rho\colon\Pi Z.\ \ell\rightarrow r\penalty 10000\ [\hskip 0.85358pt\pi\hskip 0.85358pt]$ is a left-linear constrained rule satisfying $\mathcal{V}\mathsf{ar}(\rho)\cap\mathcal{V}\mathsf{ar}(s,\varphi)=\varnothing$. We say $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ has a $\rho$-redex at position $p\in\mathcal{P}\mathsf{os}(s)$ using substitution $\gamma$ if 1. $\mathcal{D}\mathsf{om}(\gamma)=\mathcal{V}\mathsf{ar}(\ell)$, 2. $s|_{p}=\ell\gamma$, 3. $\gamma(x)\in\mathcal{V}\mathsf{al}\cup X$ for all $x\in\mathcal{V}\mathsf{ar}(\ell)\cap Z$, and 4. $\vDash_{\mathcal{M}}(\exists\vec{x}.\ \varphi)\Rightarrow(\exists\vec{z}.\ \pi\gamma)$, where $\{\hskip 0.85358pt\vec{z}\hskip 0.85358pt\}=\mathcal{V}\mathsf{ar}(\pi)\setminus\mathcal{V}\mathsf{ar}(\ell)$. Suppose $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ has a $\rho$-redex at position $p\in\mathcal{P}\mathsf{os}(s)$ using substitution $\gamma$. Then we have a most general rewrite step $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rightarrow_{\rho}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$ (or written as $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rightarrow^{p}_{\rho,\gamma}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$ with explicit $p$ and $\gamma$) where 1. $t=s[r\gamma]$, 2. $\psi=\varphi\land\pi\gamma$, 3. $\{\hskip 0.85358pt\vec{y}\hskip 0.85358pt\}=\mathcal{V}\mathsf{ar}(\psi)\setminus\mathcal{V}\mathsf{ar}(t)$, and 4. $Y=\mathcal{E}\mathsf{x}\mathcal{V}\mathsf{ar}(\rho)\cup(X\cap\mathcal{V}\mathsf{ar}(t))$. An LCTRS is a set of constrained rewrite rules that includes the set of calculation rules $\mathcal{R}_{\textsf{ca}}=\{\hskip 0.85358ptf(x_{1},\ldots,x_{n})\to y\penalty 10000\ [\hskip 0.85358pty=f(x_{1},\ldots,x_{n})\hskip 0.85358pt]\mid f\in\mathcal{F}_{\mathsf{th}}\setminus\mathcal{V}\mathsf{al}\hskip 0.85358pt\}$. However, in examples we will refrain from explicitly mentioning the calculation rules of $\mathcal{R}$. For an LCTRS $\mathcal{R}$, we have $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rightarrow_{\mathcal{R}}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$ if $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rightarrow_{\rho}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$ for some $\rho\in\mathcal{R}$. The transitive and reflexive closure $\rightarrow_{\mathcal{R}}^{*}$ of $\rightarrow_{\mathcal{R}}$ is called the most general constrained reduction.
Example 2.4.
@@
Following our formulation, the constrained rewrite rules in Example˜1.1 are presented as $\Pi\{\hskip 0.85358ptx\hskip 0.85358pt\}.\ \mathsf{sum}(x)\rightarrow 0\penalty 10000\ [\hskip 0.85358pt0\geqslant x\hskip 0.85358pt]$ and $\Pi\{\hskip 0.85358ptx\hskip 0.85358pt\}.\ \mathsf{sum}(x)\rightarrow x+\mathsf{sum}(x+-1)\penalty 10000\ [\hskip 0.85358ptx>0\hskip 0.85358pt]$. Let $s$ be the constrained term $\Pi\{\hskip 0.85358ptx\hskip 0.85358pt\}.\ \mathsf{sum}(x)\penalty 10000\ [\hskip 0.85358ptx>2\hskip 0.85358pt]$. In order to perform a rewrite step starting from $s$, we first consider a renamed variant of the second rule $\rho\colon\Pi\{\hskip 0.85358ptx^{\prime}\hskip 0.85358pt\}.\ \mathsf{sum}(x^{\prime})\rightarrow x^{\prime}+\mathsf{sum}(x^{\prime}+-1)\penalty 10000\ [\hskip 0.85358ptx^{\prime}>0\hskip 0.85358pt]$. Then $s$ admits a $\rho$-redex at position $\epsilon$ using the substitution $\gamma=\{\hskip 0.85358ptx^{\prime}\mapsto x\hskip 0.85358pt\}$, as it satisfies $\vDash_{\mathcal{M}}x>2\Rightarrow x>0$. Hence we obtain a most general rewrite step $s\to^{\epsilon}_{\rho,\gamma}\Pi\{\hskip 0.85358ptx\hskip 0.85358pt\}.\ x+\mathsf{sum}(x+-1)\penalty 10000\ [\hskip 0.85358ptx>2\land x>0\hskip 0.85358pt]$. In this case, the additional constraint $x>0$ is auxiliary. However, this addition of a constraint is necessary in general: for example, consider a rewrite rule $\rho^{\prime}:\Pi\{\hskip 0.85358ptx^{\prime},y^{\prime}\hskip 0.85358pt\}.\ \mathsf{f}(x^{\prime})\rightarrow x^{\prime}+y^{\prime}\penalty 10000\ [\hskip 0.85358pty^{\prime}\geqslant x^{\prime}\hskip 0.85358pt]$ and a most general rewrite step $\Pi\{\hskip 0.85358ptx\hskip 0.85358pt\}.\ \mathsf{f}(x)\penalty 10000\ [\hskip 0.85358ptx>2\hskip 0.85358pt]\to_{\rho^{\prime}}\Pi\{\hskip 0.85358pty^{\prime}\hskip 0.85358pt\}.\ x+y^{\prime}\penalty 10000\ [\hskip 0.85358ptx>2\land y^{\prime}>x\hskip 0.85358pt]$. Note that the extra variable $y^{\prime}$ in the rule remains uninstantiated, since the domain of the substitution is exactly $\mathcal{V}\mathsf{ar}(\ell)$.
@@ 3 Partial Reduction of Constrained Terms
Let $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ be a satisfiable existentially constrained term. Suppose that $\rho\colon\Pi Z.\ \ell\rightarrow r\penalty 10000\ [\hskip 0.85358pt\pi\hskip 0.85358pt]$ is a left-linear constrained rule satisfying $\mathcal{V}\mathsf{ar}(\rho)\cap\mathcal{V}\mathsf{ar}(s,\varphi)=\varnothing$. We say $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ has a partial $\rho$-redex at position $p\in\mathcal{P}\mathsf{os}(s)$ using substitution $\gamma$ if all of the following statements hold: 1. $\mathcal{D}\mathsf{om}(\gamma)=\mathcal{V}\mathsf{ar}(\ell)$, 2. $s|_{p}=\ell\gamma$, 3. $\gamma(x)\in\mathcal{V}\mathsf{al}\cup X$ for all $x\in\mathcal{V}\mathsf{ar}(\ell)\cap Z$, and 4. $(\exists\vec{x}.\ \varphi)\land(\exists\vec{z}.\ \pi\gamma)$ is satisfiable in $\mathcal{M}$ where $\{\hskip 0.85358pt\vec{z}\hskip 0.85358pt\}=\mathcal{V}\mathsf{ar}(\pi)\setminus\mathcal{V}\mathsf{ar}(\ell)$.
@@
We now compare partial rewrite steps with the most general rewrite steps of existentially constrained terms. First, it follows from Lemma˜3.2 that every most general rewrite step is also a partial rewrite step.
@@
As we demonstrated in Example˜3.4, partial rewrite steps may not be most general rewrite steps. Thus, the converse of Theorem˜3.5 does not hold in general. However, one can show the following characterization of the partial rewrite steps.
@@
From Lemma˜3.6, it follows that every partial rewrite step is simulated by a most general rewrite step albeit being equipped with subsumption and equivalence before and after the step.